# Settling dynamics of non-Brownian suspension of spherical and cubic particles in Stokes flow

### v1

@@ 4 Conclusion
This study presents a comprehensive investigation into the sedimentation mechanisms of cube and sphere suspensions across a broad range of solid volume fractions, from dilute to dense limits. Using the Rigid Multiblob (RMB) Method, we numerically predicted particle dynamics in a periodic domain. Our simulations uncovered fundamental aspects of particle interactions in sedimenting systems, providing insights into their effective behavior. When fluid inertia is neglected, our findings shows strong agreement with the existing empirical correlations and experimental data for spheres suspension. Particle-particle and particle-fluid interactions play a critical role in determining the macroscopic properties of sedimenting suspensions. These interactions are strongly influenced by the global particle volume concentration. We also observed that velocity fluctuations follow a $\phi^{1/2}$ scaling at low volume fractions due to microstructural effects, but at higher volume fractions, these fluctuations diminish. Moreover, we have also shown that horizontal particle velocity fluctuations are significantly higher in cube suspensions due to their susceptibility to stronger orientation-induced forces. The nature of anisotropy in velocity fluctuations diminishes monotonically from dilute to dense suspension. It is also noteworthy that some particles achieve instantaneous tendency to move against the gravity in moderate to dense regime. This is likely due to the formation of particle-rich and particle-poor regions, which induce local fluid convection, causing some particles to move upward in the particle-poor areas.
Microstructural analysis reveals that at low volume fractions, the anisotropic microstructure governs transport properties, with particles tending to orient horizontally around a test particle. This orientation effect is more pronounced for spheres than cubes, as reflected by the sharper peak in the radial distribution function (RDF) for spheres. As the volume fraction increases, interplay between many-body interactions and the microstructure control the transport properties. Furthermore, we also find that cubes are more effective in transferring momentum from the gravity to the horizontal direction, as shown by the ratio of vertical to horizontal particle velocity fluctuations. The strong transverse motions of cubes not only homogenize the suspension structure but also make particle momentum properties more isotropic compared to sphere suspensions. Additionally, in a bidisperse mixture, the settling velocities of different particle species are strongly dependent on the solid volume fraction from low to moderate values. It is observed that spheres settle faster than cubes across all ranges of $\phi$, indicating conditions that favor species segregation in these systems. Moreover, the magnitude of the horizontal and vertical velocity fluctuations increases for each species as the volume fraction rises upto a moderate limit.

### v2

@@ 4 Conclusion
This study provides a comprehensive investigation of sedimentation mechanisms in cube and sphere suspensions across a wide range of solid volume fractions, from dilute to dense regimes. Using the Rigid Multiblob (RMB) Method, we numerically predict particle dynamics within a periodic domain, uncovering key aspects of particle interactions in sedimenting systems. Our simulations align well with existing empirical correlations and experimental data for sphere suspensions when fluid inertia is neglected. Particle-particle and particle-fluid interactions play a crucial role in determining the macroscopic properties of sedimenting suspensions, with their influence strongly dependent on the global particle volume concentration. At low volume fractions, velocity fluctuations scale as $\phi^{1/2}$ due to microstructural effects but diminish at higher concentrations. Notably, horizontal velocity fluctuations are significantly higher for cubes due to stronger orientation-induced forces. Cubes also transfer momentum more effectively from gravity to the horizontal direction, as indicated by the ratio of vertical to horizontal velocity fluctuations. The anisotropy in velocity fluctuations decreases monotonically with increasing volume fraction. The strong transverse motions of cubes not only homogenize the suspension structure but also make particle momentum more isotropic compared to sphere suspensions. Additionally, in moderate to dense regimes, some particles exhibit transient upward motion against gravity. This behavior is likely driven by the formation of particle-rich and particle-poor regions, which induce local fluid convection, causing upward motion in particle-poor areas.
Microstructural analysis reveals that at low volume fractions, an anisotropic microstructure governs transport properties, with particles predominantly orienting horizontally around a test particle. This orientation effect is more pronounced for spheres than cubes, as reflected in the sharper peak of the radial distribution function (RDF) for spheres. As the volume fraction increases, many-body interactions and microstructural effects collectively control transport properties.
We also examine bidisperse suspensions composed of spherical and cubic particles, tracking the evolution of settling velocities from an initially random distribution to a steady state. In bidisperse systems, particle settling is affected by mutual interference, with settling velocities strongly dependent on solid volume fractions in the low to moderate range. Horizontal and vertical velocity fluctuations increase for each species in comparison with the monodisperse suspensions, significantly influencing suspension kinematics by enhancing anisotropy in velocity fluctuations. The presence of different particle shapes introduces inhomogeneous velocity fields due to varying drag forces, leading to differential settling speeds between species. Our results show that spheres settle faster than cubes across all volume fractions, promoting species segregation. Furthermore, increased velocity fluctuations and stronger hydrodynamic interactions in bidisperse suspensions generate locally strong return fluid fluxes, temporarily dragging more particles upward, particularly in moderate to dense regimes, highlighting the intensity of particle interactions in these conditions.